OpenStudy (anonymous):

Find two consecutive even integers such that if 6 more than the first is divided by 6 less than the second, the quotient is 2. how can i set up this problem and figure out the answer?

OpenStudy (anonymous):

let numbers be x and x + 2 x+6 --- = 2 x+2 - 6 x+6 / x - 4 = 2

OpenStudy (anonymous):

Let x and y be those two numbers. y = x+1 Also, x+6/y-6 = 2. Two equations in two variables. Solve.

OpenStudy (anonymous):

Sorry. Even. So, y = x + 2.

OpenStudy (anonymous):

solve for x x + 6 = 2x - 8 x = 14 so two numbers are 14 and 16

OpenStudy (anonymous):

ok cathy ? you can check if this is correct by going back to the questio 6 more than first = 14+6 = 20 6 less than second = 16-6 = 10 20/10 = 2

OpenStudy (anonymous):

need any more help?

OpenStudy (anonymous):

no, thanks for your help

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

plz close question by clicking on good answer